# A Differentiated Reward Method for Reinforcement Learning based Multi-Vehicle Cooperative Decision-Making Algorithms

### v1

@@ Abstract
Reinforcement learning (RL) shows great potential for optimizing multi-vehicle cooperative driving strategies through the state-action-reward feedback loop, but it still faces challenges such as low sample efficiency. This paper proposes a differentiated reward method based on steady-state transition systems, which incorporates state transition gradient information into the reward design by analyzing traffic flow characteristics, aiming to optimize action selection and policy learning in multi-vehicle cooperative decision-making. The performance of the proposed method is validated in RL algorithms such as MAPPO, MADQN, and QMIX under varying autonomous vehicle penetration. The results show that the differentiated reward method significantly accelerates training convergence and outperforms centering reward and others in terms of traffic efficiency, safety, and action rationality. Additionally, the method demonstrates strong scalability and environmental adaptability, providing a novel approach for multi-agent cooperative decision-making in complex traffic scenarios.
@@ I Introduction
As autonomous driving technology evolves towards networking and collaboration, multi-vehicle cooperative decision-making is expected to become a crucial means of enhancing traffic efficiency and road safety. Research indicates that in typical scenarios such as unsignalized intersections and highway merging zones, traditional single-vehicle decision-making systems, due to their lack of global coordination capabilities, may result in traffic efficiency loss and potential safety hazards [1]. Multi-vehicle cooperative decision-making systems hold significant research value in constructing the next generation of intelligent transportation systems.
Reinforcement Learning (RL), with its adaptive learning capabilities in dynamic environments, has gradually become one of the mainstream methods for vehicle decision-making[2, 3, 4]. Driven by deep reinforcement learning, vehicle decision systems have achieved good performance improvement in key metrics such as trajectory prediction accuracy and risk avoidance [2]. However, the application of Multi-Agent Reinforcement Learning (MARL) in autonomous driving still faces challenges such as low sample efficiency, the curse of dimensionality, and long-tail problems [5].
Multi-vehicle cooperative decision-making algorithms typically utilize vehicle speed signals, vehicle positions, and interaction events between vehicles (e.g., car-following, negotiating lane changes, collision avoidance, etc.) to design reward mechanisms. These signals help guide the vehicles to make reasonable decisions. Therefore, the design of the reward function is of crucial importance [6, 7]. Specifically, from a mesoscopic traffic flow perspective, the state of vehicles is in most time stable and changes gradually over time. This can lead to reinforcement learning algorithms failing to distinguish between actions due to errors. In this paper, we proposed a differentiated reward method based on a steady-state transition system. Experimental results demonstrate that the differentiated reward method significantly accelerates the training convergence speed of reinforcement learning and exhibits more rational behavior in action selection.
@@ II Related works
Multi-Vehicle Cooperative Decision-Making Based on Reinforcement Learning: In recent years, the problem of multi-vehicle cooperative decision-making in dynamic traffic environments has garnered significant attention. Rule-based or optimization-based methods often struggle to scale and adapt in complex traffic scenarios. Increasingly mature reinforcement learning algorithms such as MADQN, MADDPG, MAPPO, and Qmix enable agents to learn optimal strategies through interaction with the environment [4]. These algorithms have significantly improved the quality of multi-vehicle interaction decision-making in scenarios such as unsignalized intersections [8, 9, 10, 11], highway ramps [12, 13, 14], and mixed scenarios [15, 16].
Despite the success of existing multi-vehicle cooperative decision-making methods in certain fixed scenarios, designing efficient reinforcement learning algorithms that account for the dynamic characteristics of traffic flow remains an open research challenge in complex, continuous traffic flow environments.
Reward Function Design in RL: In reinforcement learning, the reward function is crucial for guiding the learning process of agents. Designing an appropriate reward function can significantly enhance the performance of algorithms [17, 18], especially in complex multi-vehicle cooperative decision-making problems, where it is essential to incorporate problem-specific features into the reward design.
Reward shaping is a commonly used technique in reinforcement learning research. For instance, [19] and [20] introduced a bi-level optimization mechanism to improve data efficiency, enabling effective learning of reward functions and policies even with limited human feedback.
Building on the aforementioned research, this study introduces a differential reward mechanism to enhance multi-vehicle cooperative decision-making capabilities. By leveraging reward differentiation, we propose a theoretically grounded approach to improve policy learning and multi-agent coordination. Specifically, in continuous traffic flow environments, this method effectively optimizes state transitions and stabilizes the reward distribution among agents, thereby enhancing decision-making performance.
@@ III-C Observation/State Space
For the agent vehicle $i$, its own state vector is defined as:
$s_{i}^{\text{self}}=[p_{i}^{\text{lon}},p_{i}^{\text{lat}},v_{i},\tau_{i},g_{i},d_{i}^{\text{left}},d_{i}^{\text{front}},d_{i}^{\text{right}}]$
@@
For the set of surrounding vehicles $\text{nbr}^{i}=\{j\mid\|p_{i}-p_{j}\|_{2}\leq R\}$ (where $R$ is the perception radius), construct the relative state matrix:
$M_{i}^{\text{nbr}}=\underset{j\in\text{nbr}^{i}}{\text{concat}}\left[\Delta p_{ij}^{\text{lon}},\Delta p_{ij}^{\text{lat}},\Delta v_{ij},\Delta\tau_{ij},\Delta g_{ij}\right],$
where $\Delta p_{ij}^{\text{lon}}=p_{j}^{\text{lon}}-p_{i}^{\text{lon}}$ and $\Delta p_{ij}^{\text{lat}}=p_{j}^{\text{lat}}-p_{i}^{\text{lat}}$ represent the relative positional relationships, $\Delta v_{ij}=v_{j}-v_{i}$ denotes the speed difference, $\Delta\tau_{ij}=\tau_{j}-\tau_{i}$ reflects the difference in vehicle types, and $\Delta g_{ij}=\|g_{j}-g_{i}\|_{2}$ is the Euclidean distance representing the difference in driving goals. When the number of surrounding vehicles $n<N_{max}$, the missing rows are padded with zero vectors.
@@
$o_{i}=\text{concat}\left[s_{i}^{\text{self}},M_{i}^{\text{nbr}}\right]$
@@ III-D Action space
$\mathcal{A}_{i}^{\text{lon}}=\{a^{\text{acc}},a^{\text{keep}},a^{\text{dec}}\}$
@@
$\dot{v}_{i}=\text{clip}(v_{i}+a^{\text{lon}}\cdot\Delta t,0,v_{\text{max}})$
@@
$\mathcal{A}_{i}^{\text{lat}}=\{a^{\text{left}},a^{\text{hold}},a^{\text{right}}\}$
@@
$L_{i}^{t+1}=\begin{cases}\max(L_{i}^{t}-1,1)&\text{if }a^{\text{lat}}=a^{\text{left}}\\ L_{i}^{t}&\text{if }a^{\text{lat}}=a^{\text{hold}}\\ \min(L_{i}^{t}+1,N_{lane})&\text{if }a^{\text{lat}}=a^{\text{right}}\end{cases}$
@@
$\mathcal{A}_{i}=\mathcal{A}_{i}^{\text{lon}}\times\mathcal{A}_{i}^{\text{lat}}$
@@ IV Methodology
IV-A Introduction of Reward Centering
@@
where $r(\pi)$ is the state-independent average reward obtained by policy $\pi$, and $\tilde{h}_{\pi}(s)$ is the differential value of state $s$. For ergodic Markov decision processes, these two terms can be defined as Equation 2.
@@
$\tilde{h}_{\pi}(s)$ $\doteq\mathbb{E}\left[\sum_{k=1}^{\infty}\big{(}R_{t+k}-r(\pi)\big{)}\mid S_{t}=s,A_{t:\infty}\sim\pi\right],\vspace{-0.5mm}$
here, $e_{\pi}^{\gamma}(s)$ represents an error term that approaches zero as the discount factor approaches 1. To distinguish the speed $v$, we do not use the common value function notation in reinforcement learning research but instead use $h$ to denote the value function.
@@
IV-B Reward Differentiation and Its Connection to Reward Centering
Next, we formulate differentiated reward and analyze its intrinsic connection to reward centering.
Define a Markov chain $\mathcal{S}=\{S(t):t=0,1,2,\ldots\}$, whose state space is $\Re^{\ell}$, and $\mathcal{S}$ evolves according to a nonlinear state-space model:
$S(t+1)=a(S(t),N(t+1)),\quad t\geq 0,$ (4)
Under these assumptions, for all $t\geq 0$, $S(t+1)$ is a continuous function of the initial condition $S(0)=s_{0}$.
@@ IV-B Reward Differentiation and Its Connection to Reward Centering
Assume that the value function $h^{\gamma}$ and all its possible approximations $\{h_{\pi(\theta)}^{\gamma}:\theta\in\Re^{d}\}$ are continuously differentiable as functions of the state $s$, i.e., for each $\theta\in\Re^{d}$, $h^{\gamma},\,h_{\pi(\theta)}^{\gamma}\in C^{1}$. Based on the linear parameterization in (6), we obtain the following form of the differential value function:
@@
$\theta^{*}$ $=\operatorname*{arg\,min}_{\theta}\|\nabla h_{\pi(\theta)}^{\gamma}-\nabla h^{\gamma}\|^{2}_{\pi}$ (9)
Paper [22] provides a detailed convergence proof of the problem represented by Equation (9). Clearly, in general reinforcement learning algorithms, directly using the differential reward function is essentially equivalent to solving the problem posed by Equation (9).
In steady-state traffic flow, if the following conditions hold:
  • Condition 1: The $\nabla h^{\gamma}$ is linearly correlated with the state transition direction $\mathbb{E}[S_{t+1}|S_{t}=s]$;
  • Condition 2: The reward differential weight $\lambda$ and the centered mean $r(\pi)$ satisfy $\lambda\cdot\nabla h^{\gamma}\propto\tilde{R}_{t}$,
then reward differentiation and reward centering are equivalent in the value function update, i.e.,
$\mathbb{E}_{\pi}\left[\tilde{R}_{t}+\lambda\cdot r_{p}^{i}\right]\approx\mathbb{E}_{\pi}\left[\tilde{R}_{t}\right].$
At this point, both methods adjust the distribution of reward signals to make the value function focus more on the relative differences between states.
The effectiveness of reward centering has been demonstrated in paper [19]. Therefore, in reinforcement learning problems with steady-state transitions, the reward differentiation method is expected to improve performance.
IV-C Differentiated Reward Implementation in Vehicle Decision-Making
In the field of vehicle decision-making (multi-vehicle collaborative decision-making), the reward function commonly used can be expressed as Equation (10). [23, 24, 25]
@@ IV-C Differentiated Reward Implementation in Vehicle Decision-Making
$=\frac{1}{N}\left(w_{1}{\sum_{i=1}^{N}\frac{v_{i}}{v_{\max}}}+w_{2}N_{\text{sat}}+w_{3}N_{\text{col }}+w_{4}N_{LC}\right)$
@@
Due to the variable number of agents in the setup, this paper cannot directly use the sum of the local rewards of all agents as the reward function. Instead, an index needs to be designed to objectively evaluate the overall traffic quality within the coordinating zone. To this end, we design the following external reward function:
$r_{\text{env}}=\frac{1}{\left|\mathcal{N}_{\text{CAV}}\right|}\textstyle\sum_{i}\left(\omega_{1}r_{a}^{i}+\omega_{2}r_{p}^{i}\right)+\omega_{3}r_{\text{flow}}+\omega_{4}r_{\text{safe}}$ (11)
@@
$r_{a}=\begin{cases}1&\text{ if accelerating or keeping highspeed,}\\ 0&\text{ otherwise.}\end{cases}$
We design a potential field based on the vehicle’s longitudinal position and the lane it occupies, to evaluate the value of the current position of vehicle $i$ relative to its target.
$f_{p}^{i}\left(x,y\right)=\frac{e^{-\frac{(l-x)^{2}}{2\sigma^{2}}}}{\zeta\left|y_{tar}^{i}-y\right|+1}$ (12)
@@
$r_{p}^{i}=\mathbf{v}^{i}\cdot\nabla f_{p}^{i}\left(x,y\right)$ (13)
@@
$r_{p}^{i}=\left[v_{x}^{i}\left(l-x\right)+\frac{\zeta v_{y}^{i}\text{sign}\left(y-y_{tar}^{i}\right)}{\zeta\left|y_{tar}^{i}-y\right|+1}\right]\cdot f_{p}^{i}\left(x,y\right)$ (14)
@@
$r_{\text{flow}}=\frac{1}{\left|\mathcal{N}\right|}{\textstyle\sum_{i}}\frac{v^{i}}{v_{max}}$
For the safety indicator, since accidents are rare events and their occurrence typically has a significant impact on overall traffic, we use summation rather than averaging.
$r_{\text{safe}}={\textstyle\sum_{i}}\mathbb{I}(i)$

### v2

@@ Abstract
Deep reinforcement learning (DRL) shows great potential for optimizing multi-vehicle cooperative driving strategies through the state-action-reward feedback loop, but it still faces challenges such as low sample efficiency. This paper proposed a differentiated reward method based on steady-state transition systems, which incorporates state transition gradient information into the reward design by analysing traffic flow characteristics, aiming to optimize action selection and boost policy learning in multi-vehicle cooperative decision-making. The performance of the proposed method is validated in RL algorithms such as MAPPO, MADQN, and QMIX under varying connected and automated vehicle (CAV) penetration rate. Results show that the differentiated reward method significantly accelerates training convergence and outperforms centering reward and other common reward shaping method in terms of traffic efficiency, safety, and action rationality. Additionally, the method demonstrates strong scalability and environmental adaptability, providing a novel approach for multi-agent cooperative decision-making in complex traffic scenarios. Code is available at https://github.com/leoPub/diff_rew.
@@ I Introduction
With continuous advancements in perception technologies and local path planning for single-vehicle autonomous driving systems, coupled with rapid development of V2X communication and computing platforms, multi-vehicle collaborative decision-making has demonstrated significant potential as a key approach to enhance traffic efficiency and road safety. Research indicates that in typical scenarios such as unsignalized intersections and highway merging zones, single-vehicle decision-making systems may lead to traffic efficiency degradation and safety risks due to insufficient global coordination capabilities [1]. The vehicle cooperative decision-making systems hold crucial research value for constructing next-generation intelligent transportation system.
Deep reinforcement learning (DRL), with its self-adaptive learning capabilities in dynamic environments, has emerged as one of the mainstream methods for vehicle decision-making [2, 3, 4]. DRL-based vehicle decision-making algorithms have demonstrated significant performance improvements in critical metrics such as trajectory prediction accuracy and risk avoidance [2]. However, the application of Multi-Agent Reinforcement Learning (MARL) in multi-vehicle decision making (MVDM) still faces persistent challenges including low sample efficiency, the curse of dimensionality, and long-tail problems [5].
In MVDM researches, reward functions are typically designed based on vehicle positions, velocities, and inter-vehicle interaction events (e.g., car-following, lane-changing negotiation, collision avoidance). Reward functions guide vehicles to make rational decisions, rendering its great importance in MVDM algorithms [6, 7]. Specifically, from a mesoscopic traffic flow perspective, vehicle states remain stable and evolve gradually over time in most scenarios. Under typical decision-making frequencies ($\geq 10\text{Hz}$), reinforcement learning algorithms may fail to distinguish between different actions in adjacent states due to data noises. This paper proposes a differentiated reward approach based on a steady-state transition system. Experimental results demonstrate that the proposed method significantly accelerated the training convergence speed of reinforcement learning and exhibited more rational behavior in action selection.
@@ II Related works
DRL-based Multi-Vehicle Cooperative Decision-Making: In recent years, the problem of multi-vehicle cooperative decision-making in dynamic traffic environments has garnered significant attention. Rule-based or optimization-based methods often struggle to scale and adapt in complex traffic scenarios. Increasingly mature reinforcement learning algorithms such as MADQN, MADDPG, MAPPO, and Qmix enable agents to learn optimal strategies through interaction with the environment [4]. These algorithms have significantly improved the quality of multi-vehicle interaction decision-making in scenarios such as unsignalized intersections [8, 9, 10, 11], highway ramps [12, 13, 14], and mixed scenarios [15, 16]. Zhuang H et al. proposed a distributed multi-agent proximal policy optimization (MAPPO) method with an attention-enhanced mechanism (Attn-MAPPO) for joint collision avoidance and efficient intersection traversal, demonstrating superior performance over heuristic rule-based models [][8]. Chen D et al. formulated the mixed-traffic highway ramp merging problem as a MARL task, employing curriculum learning to effectively train agents for challenging driving maneuvers [][13]. Current research demonstrates that reinforcement learning-based multi-vehicle cooperative decision-making algorithms have achieved human-like or even superhuman driving capabilities. However, enhancing training sample efficiency and stability remains an open problem.
Reward Function Design in DRL: In reinforcement learning, the reward function plays a pivotal role in guiding the learning process of agents. A properly designed reward function can significantly enhance algorithm performance [17, 18], particularly in complex multi-vehicle collaborative decision-making problems, where incorporating problem-specific characteristics into reward design is particularly of critical importance.
Reward shaping is a commonly employed technique in reinforcement learning (RL) research that incorporates domain knowledge into RL frameworks to accelerate the discovery of optimal policies[][19]. R. D. et al. proposed a straightforward framework for integrating shaped rewards into RL systems and accompanied this with analytical tools to demonstrate that specific reward shaping choices can significantly enhance sample efficiency [][20]. A generally applicable reward centering approach for RL is introduced in [][21], which substantially improves the performance of RL algorithms under standard discount factors, however struggles to handle scenarios where reward functions contain substantial underlying offsets.
Based on the aforementioned research, this study proposes a differentiated reward mechanism to enhance multi-vehicle cooperative decision-making capabilities in RL. We establish a theoretical framework for reward- differentiating reinforcement learning and demonstrate its application in multi-vehicle coordination through concrete case studies. In continuous traffic flow environments, the proposed method effectively optimizes inter-agent reward allocation during steady-state transitions, thereby enhancing sample efficiency, accelerating convergence speed and stability in algorithm training.
@@ III-C Observation/State Space
For the vehicle $i$, its own state vector is defined as
$s_{i}^{\text{self}}=[p_{i}^{\text{lon}},p_{i}^{\text{lat}},v_{i},\tau_{i},g_{i},d_{i}^{\text{left}},d_{i}^{\text{front}},d_{i}^{\text{right}}],$ (1)
@@
For the set of surrounding vehicles $\text{nbr}^{i}=\{j\mid\|p_{i}-p_{j}\|_{2}\leq R\}$ (where $R$ is the perception radius), we construct the following relative state matrix:
$M_{i}^{\text{nbr}}=\underset{j\in\text{nbr}^{i}}{\text{concat}}\left[\Delta p_{ij}^{\text{lon}},\Delta p_{ij}^{\text{lat}},\Delta v_{ij},\Delta\tau_{ij},\Delta g_{ij}\right],$ (2)
where $\Delta p_{ij}^{\text{lon}}=p_{j}^{\text{lon}}-p_{i}^{\text{lon}}$ and $\Delta p_{ij}^{\text{lat}}=p_{j}^{\text{lat}}-p_{i}^{\text{lat}}$ represent the relative positional relationship, $\Delta v_{ij}=v_{j}-v_{i}$ denotes the velocity difference, $\Delta\tau_{ij}=\tau_{j}-\tau_{i}$ reflects the vehicle type difference, and $\Delta g_{ij}=\|g_{j}-g_{i}\|_{2}$ represents the driving goal difference in terms of Euclidean distance. When the number of surrounding vehicles $n<N_{max}$, the missing rows are padded with zero vectors.
@@
$o_{i}=\text{concat}\left[s_{i}^{\text{self}},M_{i}^{\text{nbr}}\right].$ (3)
@@ III-D Action space
$\mathcal{A}_{i}^{\text{lon}}=\{a^{\text{acc}},a^{\text{keep}},a^{\text{dec}}\},$ (4)
@@
$\dot{v}_{i}=\text{clip}(v_{i}+a^{\text{lon}}\cdot\Delta t,0,v_{\text{max}}),$ (5)
@@
$\mathcal{A}_{i}^{\text{lat}}=\{a^{\text{left}},a^{\text{hold}},a^{\text{right}}\},$ (6)
@@
$L_{i}^{t+1}=\begin{cases}\max(L_{i}^{t}-1,1)&\text{if }a^{\text{lat}}=a^{\text{left}}\\ L_{i}^{t}&\text{if }a^{\text{lat}}=a^{\text{hold}}\\ \min(L_{i}^{t}+1,N_{lane})&\text{if }a^{\text{lat}}=a^{\text{right}}\end{cases},$ (7)
@@
$\mathcal{A}_{i}=\mathcal{A}_{i}^{\text{lon}}\times\mathcal{A}_{i}^{\text{lat}}.$ (8)
@@ IV Methodology
IV-A Method of Reward Centering
@@
where $r(\pi)$ is the state-independent average reward obtained by policy $\pi$, and $\tilde{h}_{\pi}(s)$ is the differential value of state $s$. For ergodic Markov decision processes, these two terms can be defined as Equation (10).
@@
$\tilde{h}_{\pi}(s)$ $\doteq\mathbb{E}\left[\sum_{k=1}^{\infty}\big{(}R_{t+k}-r(\pi)\big{)}\mid S_{t}=s,A_{t:\infty}\sim\pi\right].\vspace{-0.5mm}$
Additionally, $e_{\pi}^{\gamma}(s)$ represents an error term that approaches zero as the discount factor approaches 1. To distinguish the speed $v$, we do not use the common value function notation in reinforcement learning research but instead use $h$ to denote the value function.
@@
IV-B Differentiated Reward Formulation and Analysis
We define a Markov chain $\mathcal{S}=\{S(t):t=0,1,2,\ldots\}$, whose state space is $\Re^{\ell}$, and $\mathcal{S}$ evolves according to a nonlinear state-space model:
$S(t+1)=\mathcal{T}(S(t),N(t+1)),\quad t\geq 0,$ (12)
where $N$ is an m-dimensional disturbance sequence modeling the nonlinear state transition, composed of independent and identically distributed random variables, $\mathcal{T}$ is a continuous mapping. Under this assumption, for all $t\geq 0$, $S(t+1)$ is a continuous function of the initial condition $S(0)=s_{0}$.
@@ IV-B Differentiated Reward Formulation and Analysis
Assume that the value function $h^{\gamma}$ and all its possible approximations $\{h_{\pi(\theta)}^{\gamma}:\theta\in\Re^{d}\}$ are continuously differentiable as functions of the state $s$, i.e., for each $\theta\in\Re^{d}$, $h^{\gamma},\,h_{\pi(\theta)}^{\gamma}\in C^{1}$. Based on the linear parameterization in Equation (14), we obtain the following form of the differential value function:
@@
$\theta^{*}$ $=\operatorname*{arg\,min}_{\theta}\|\nabla h_{\pi(\theta)}^{\gamma}-\nabla h^{\gamma}\|^{2}_{\pi}.$ (17)
Evidently, in conventional reinforcement learning algorithms, directly utilizing the derivative of the reward function is essentially equivalent to solving the problem posed by Equation (17).
It can be observed that both reward centering and the reward differentiating method exhibit translation invariance to shifts in the reward function (removing the mean or computing gradients can eliminate the effects of constant offsets). When $\gamma\rightarrow 1$ ,the long-term fluctuations of ${h}_{\pi\left(s\right)}^{\gamma}\left(s\right)$ under steady-state conditions in reward centering are averaged out,causing the value function to be predominantly governed by $r\left(\pi\right)/\left({1-\gamma}\right)$ The centered value function satisfies Bellman equation ${\widetilde{h}}_{\pi}^{\gamma}\left(s\right)=\mathbb{E}\left[{R-r\left(\pi\right)+\gamma{\widetilde{h}}_{\pi}^{\gamma}\left({s}^{\prime}\right)\mid s}\right]$ ,while the gradient objective $\nabla{h}_{x}^{\gamma}\left(s\right)$ in the reward differentiating method aligns with the solution of the steady-state Bellman equation.
Furthermore, the reward differentiating method exhibits superior performance over reward centering when addressing scenarios involving global constant offsets or state-dependent systematic biases in the reward function. Consider a reward function with a global constant offset $c$ Reward centering requires precise estimation of $r\left(\pi\right)=c+\mathbb{E}\left[{R}_{r}\right]$ ,where significant estimation errors under large offsets may introduce biases in value function estimation. For state-dependent systematic biases (e.g., $R\left(s\right)={R}_{\text{true }}\left(s\right)+c\left(s\right)$ ,where $c\left(s\right)$ denotes a state-dependent offset function),reward centering necessitates estimating the average offset $\mathbb{E}\left[{c\left(s\right)}\right]$ for each state,which becomes computationally intractable in complex environments. In contrast,the gradient $\nabla{h}^{\gamma}\left(s\right)$ is driven by difference of rewards between adjacent states. Even when state-dependent offsets $c\left(s\right)$ exist,as long as the offset function varies smoothly across local state transitions (e.g., $c\left({s}^{\prime}\right)-c\left(s\right)$ remains small),the gradient updates can still approximate the true reward difference ${R}_{\text{true }}\left({s}^{\prime}\right)-{R}_{\text{true }}\left(s\right)$
IV-C Differentiated Reward Implementation in MVDM
In MVDM, the reward function commonly used can be expressed as Equation (18). [23, 24, 25]
@@ IV-C Differentiated Reward Implementation in MVDM
$=\frac{1}{N}\left(w_{1}{\sum_{i=1}^{N}\frac{v_{i}}{v_{\max}}}+w_{2}N_{\text{sat}}+w_{3}N_{\text{col }}+w_{4}N_{LC}\right),$
@@
Due to the variable number of agents in the setup, this paper cannot directly use the sum of the local rewards of all agents as the reward function. Instead, an index needs to be designed to objectively evaluate the overall traffic quality within the coordinating zone. To this end, we design the following reward function:
$r_{\text{env}}=\frac{1}{\left|\mathcal{N}_{\text{CAV}}\right|}\textstyle\sum_{i}\left(\omega_{1}r_{a}^{i}+\omega_{2}r_{p}^{i}\right)+\omega_{3}r_{\text{flow}}+\omega_{4}r_{\text{safe}},$ (19)
@@
$r_{a}=\begin{cases}1&\text{ if accelerating or keeping highspeed,}\\ 0&\text{ otherwise.}\end{cases}$ (20)
We design a potential field based on the vehicle $i$’s longitudinal position and the lane it occupies, to evaluate the value of the current position of vehicle $i$ relative to its target.
$f_{p}^{i}\left(x,y\right)=\frac{e^{-\frac{(l-x)^{2}}{2\sigma^{2}}}}{\zeta\left|y_{tar}^{i}-y\right|+1},$ (21)
@@
$r_{p}^{i}=\mathbf{v}^{i}\cdot\nabla f_{p}^{i}\left(x,y\right),$ (22)
@@
$r_{p}^{i}=\left[v_{x}^{i}\left(l-x\right)+\frac{\zeta v_{y}^{i}\text{sign}\left(y-y_{tar}^{i}\right)}{\zeta\left|y_{tar}^{i}-y\right|+1}\right]\cdot f_{p}^{i}\left(x,y\right),$ (23)
@@
$r_{\text{flow}}=\frac{1}{\left|\mathcal{N}\right|}{\textstyle\sum_{i}}\frac{v^{i}}{v_{max}}.$ (24)
For the safety indicator, since accidents are rare events and their occurrence typically has a significant impact on overall traffic, we use summation rather than averaging, that is
$r_{\text{safe}}={\textstyle\sum_{i}}\mathbb{I}(i),$ (25)